TWO DIMENSIONAL COHERENT DOUBLE RESONANCE ELECTRONIC SPECTROSCOPY

\underline{PETER C. CHEN}, Spelman College, 350 Spelman Lane SW, Atlanta, GA 30314.

Coherent Double Resonance Electronic Spectroscopy can be used to create 2D spectra that sort peaks by rotational and vibrational quantum numbers, vibrational sequence, rotational selection rule, and species. This talk will include a brief introduction to the technique and examples of these 2D spectra.